# On the non-Markovian quantum control dynamics

### v2

@@ I Introduction
Quantum control has garnered much attention due to its potential applications in quantum optics [1, 2], quantum information processing [3, 4, 5, 6], quantum engineering [7] and others [8, 9, 10, 11, 12, 13]. Control methods for quantum systems can be categorized into the open-loop control and closed-loop control, similar to that in classical systems [1]. In open-loop control, designed or iteratively optimized control pulses are utilized to produce required gate operations in quantum computations [14, 15, 16, 17]. When a control field is applied upon an atom, the atom can be excited, leading to the generation of single or multiple photons for quantum networking [18]. On the other hand, closed-loop quantum control can be realized through coherent feedback or measurement feedback methods. For instance, when an atom or cavity quantum electrodynamics (cavity-QED) system is coupled with a waveguide, a coherent feedback channel can be established using photons transmitted in the waveguide [19]. Subsequently, the atomic dynamics and photonic states can be modulated by tuning the parameters of the coherent feedback loop [20, 21, 19, 22, 23]. This form of coherent feedback dynamics can be represented as a linear control system with delays determined by the length of the feedback loop [21, 19]. Then the quantum state dynamics can be interpreted in terms of the stability of a linear control system with delays [22, 23, 24].
@@
In addition to quantum coherent feedback, quantum measurement feedback is another commonly employed method for feedback control in quantum systems [1]. This approach involves designing feedback control based on the measurement outcomes of the quantum system [25, 24]. The measurement of quantum states can be categorized into quantum non-demolition (QND) measurement and non-QND measurement, depending on whether the measurement operator commutes with the quantum system’s Hamiltonian [1, 24]. These measurement methods have an impact on the steady quantum states [26, 27, 28, 29]. In the realm of control theory, quantum measurements can be represented using stochastic equations due to measurement and detection noise [30, 31, 32], distinguishing if from feedback control without measurement [33, 30]. By employing these measurement techniques and feedback control, quantum measurement feedback can influence the evolution of quantum states and facilitate the generation of desired quantum states. Consequently, quantum measurement feedback control holds significant promise for various applications in open quantum systems where the quantum states are affected by the environment [1]. One notable application is the use of measurement-based quantum feedback control in the quantum error correction (QEC) to rectify error bits in quantum computations [34, 35, 36]. Additionally, it can help preserve the coherence of a quantum state when the quantum system interacts with its environment[37].
@@
In the realm of open quantum system control, it is typically assumed that the environmental evolution timescale is considerably shorter than the atomic system, and further can be assumed as static [38]. This assumption allows for the modeling of the interaction between the quantum system and the environment using a master equation with a static decaying rate, which is a widely employed Markovian approximation method for studying quantum dynamics in open systems. However, in numerous scenarios, this Markovian approximation proves inadequate for comprehensively analyzing the dynamics of open quantum systems [38, 39]. For instance, in the experimental implementation utilizing nuclear magnetic resonance (NMR) [40], where the NMR qubit interacts with a non-Markovian environment characterized by a randomized configuration of modulated radio-frequency fields. In such setups, the interaction between the NMR and the environment can lead to information backflow from the environment to the qubit, thus the decoherence of the qubit can be nonmonotonic, which is different from the monotonic decoherence of a qubit in the Markovian environment [41]. Moreover, experimental evidence of information backflow induced by non-Markovianity has also been observed in optical systems [42]. Given that traditional quantum control strategies relying on Markovian approximation are not directly applicable in these instances, it is required to explore the open-loop and closed-loop quantum control dynamics for the non-Markovian open quantum systems. Additionally, it is worth noting that the traditional Morkovian scenario can be encompassed within the broader framework of non-Markovian settings as a simplified special case.
@@
The interaction between the quantum system and its environment, characterized as non-Markovian, can be effectively represented by the stochastic Schrödinger equation [43]. This approach incorporates the influence of the environment on quantum states through a complex-valued stochastic process [44, 43]. Alternatively, the non-Markovian dynamics can be described using a master equation featuring time-varying Lindblad components [45, 46], which also represents the memory effect inherent in non-Markovian dynamics, distinguish it from the Markovian master equation [47]. Consequently, various non-Markovian quantum control techniques have been developed, expanding upon the foundation laid by Markovian quantum control methods [48, 49, 50].
@@
In this paper, we utilize the commonly employed cavity-QED system depicted in Fig. 1 as an example to investigate the dynamics of non-Marovkovian quantum control in a novel manner. We approach the evolution of parameters in quantum non-Markovian control as a nonlinear control problem, resulting in the introduction of time-varying parameters in the linear evolution of quantum states due to non-Markovianity. Initially, we analyze the non-Markovian interactions between the atom and the environment from a control perspective, and for the first time analyze the transition from non-Markovian transient dynamics to steady Markovian dynamics using stability theory in nonlinear control. Subsequently, we describe the evolution of atomic and photonic states using linear time varying (LTV) control equations, where external drives or the measurement feedback can influence the equation dynamics. Lastly, we extend our analysis to encompass non-Markovian dynamics in systems where atoms in multiple coupled cavities interact with a non-Markovian environment. Then the quantum states can be categorized into the stable and unstable subspaces based on the application of measurement feedback controls.
@@ II Non-Markovian quantum control for open systems
Figure 1: Quantum control based on the Markovian interactions between a multi-level atom and a cavity, as well as the non-Markovian interactions between the atom and the environment modeled as a collection of oscillators.
As illustrated in Fig. 1, a resonant cavity is constructed between two mirrors, one multi-level atom with the energy levels represented with the state vectors $|0\rangle,|1\rangle,\cdots,|N-1\rangle$ is coupled with the cavity. The multi-level atom is simultaneously coupled with the environment modeled as a group of oscillators represented with the yellow circles. The cavity can be driven by the field represented with the red arrow, and can be detected or measured via its output channel, which is represented with the blue arrow. Using the measurement information, the feedback operator $G$ can be designed and applied upon the quantum system, thus the closed-loop control with measurement feedback can be realized.
@@
where $\omega_{c}$ is the resonant frequency of the cavity with the annihilation (creation) operator $a$ ($a^{{\dagger}}$), the second component represents a group of oscillators in the environment modeled by the annihilation (creation) operator $b_{\omega}$ ($b_{\omega}^{{\dagger}}$) with different frequencies $\omega$, the third component is the atom Hamiltonian with $\omega_{n}$ represents the energy of the $n$-th level represented with the multiplication of the vector $|n\rangle$ and $\langle n|$, $H_{t}$ is composed with the interaction Hamiltonian between the atom and cavity, as well as the interaction Hamiltonian between the atom and environment. According to Ref. [51], the environment can be represented with the oscillator $\left|z_{\omega}\right\rangle=\exp\left(z_{\omega}b_{\omega}^{{\dagger}}\right)\left|0\right\rangle$.
The analysis on the dynamics among the atom, cavity and environment in this paper, as well as the meaning of the $H_{t}$ in Eq. (1), is based on the following assumptions.
Assumption 1.
The interaction between the cavity and atom is Markovian, and the interaction between the atom and environment is initially non-Markovian.
Assumption 2.
([52, 53]) The transition frequency between two neighborhood energy levels are identical, i.e., $\omega_{n}-\omega_{n-1}\equiv\omega_{a}$ for arbitrary $n$.
@@
$\bar{H}_{t}=$ $\sum_{n=1}^{N-1}g_{n}\left(e^{-i\delta t}\sigma^{-}_{n}a^{{\dagger}}+e^{i\delta t}\sigma^{+}_{n}a\right)+\sum_{n=1}^{N-1}\sum_{\omega}\chi_{\omega}^{(n)}\left(\sigma^{-}_{n}b_{\omega}^{{\dagger}}+\sigma^{+}_{n}b_{\omega}\right),$ (2)
where $g_{n}$ represents the coupling strengths between the cavity and different atom energy levels, $\delta=\omega_{a}-\omega_{c}$ represents the detuning between the cavity resonant frequency and the transition frequency of two neighborhood atom energy levels, the lowering operator $\sigma^{-}_{n}=|n-1\rangle\langle n|$, and the rising operator $\sigma^{+}_{n}=|n\rangle\langle n-1|$. The first part on RHS represents the interaction between the atom and cavity, $\sigma^{-}_{n}a^{{\dagger}}$ represents that when the atom decays from the $n$-th energy level to the $(n-1)$-th energy level, it can emit one photon into the cavity. The following Hermitian conjugate $\sigma^{+}_{n}a$ represents that the atom can absorb one photon from the cavity and be excited to a higher energy level. The second part at RHS similarly represents the interaction between the multi-level atom and environment, and the coupling strength between the $n$-th energy level and the environmental oscillator with the frequency $\omega$ is $\chi_{\omega}^{(n)}$. $\sigma^{-}_{n}b_{\omega}^{{\dagger}}$ represents that an emitted filed generated by the atom’s decay from the $n$-th to $(n-1)$-th energy level can be absorbed by the environment, similar for the following Hermitian conjugate.
The overall coupling between the multi-level atom and the environment can be represented with the operator [55]
$L=\sum_{n=1}^{N-1}\kappa_{n}|0\rangle\langle n|,$ (3)
where $\kappa_{n}$ represents the decaying of the atomic state $|n\rangle$ to the environment. More details are introduced in Appendix A.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}|\psi(t)\rangle=-i\bar{H}|\psi(t)\rangle+L|\psi(t)\rangle z_{t}-L^{{\dagger}}\int_{0}^{t}\alpha(t,s)\frac{\delta|\psi(t)\rangle}{\delta z_{s}}\mathrm{d}s,$ (4)
where $\bar{H}$ represents the interaction component between the atom and cavity in Eq. (2), $z_{t}$ is a complex stochastic process with $E(z_{t}^{*}z_{s})=\alpha(t,s)$, $E(z_{t}z_{s})=0$, and
@@
Then the density matrix representation of quantum states is governed by the following master equation according to the derivation in Appendix B,
$\dot{\rho}(t)=$ $-i[\bar{H},\rho(t)]$ (6)
$+\left(\int_{0}^{t}\alpha^{*}(t,s)f^{*}(t,s)\mathrm{d}s+\int_{0}^{t}\alpha(t,s)f(t,s)\mathrm{d}s\right)L\rho(t)L^{{\dagger}}$
$-\int_{0}^{t}\alpha(t,s)f(t,s)\mathrm{d}sL^{{\dagger}}L\rho(t)-\int_{0}^{t}\alpha^{*}(t,s)f^{*}(t,s)\mathrm{d}s\rho(t)L^{{\dagger}}L,$
where we denote
$F(t)=\int_{0}^{t}\alpha(t,s)f(t,s)\mathrm{d}s,$ (7)
and the integration in Eq. (6) represents the memory effect of the non-Markovian process based on the interaction between the atom and the environment. We can distinguish the Markovian and non-Markvoian quantum dynamics in the open systems according to the following definition.
Definition 1.
The quantum dynamics described by the master equation (6) is non-Markovian when $F(t)$ is a time-dependent integration with a non-Markovian kernel, and is Markovian when $F(t)$ equals a constant.
@@
When $\gamma\rightarrow\infty$, the interaction between the system and the environment is Markovian and $\alpha(t,s)=\delta(t-s)$ [44]. Then
$F(t)=\int_{0}^{t}\delta(t-s)f(t,s)\mathrm{d}s=\frac{1}{2}f(t,t)\triangleq\frac{1}{2}\chi,$ (8)
is a constant and the dynamics reduces to the case in open quantum systems with Markovian interactions between the atom and environment.
@@ II-A Single multi-level atom coupled with the cavity
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle=-ig_{n}e^{i\delta t}\langle\sigma^{+}_{n}a\rangle+ig_{n}e^{-i\delta t}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+ig_{n+1}e^{i\delta t}\langle\sigma^{+}_{n+1}a\rangle-ig_{n+1}e^{-i\delta t}\langle\sigma^{-}_{n+1}a^{{\dagger}}\rangle$
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\left(F(t)+F^{*}(t)\right)|\kappa_{n}|^{2}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle,$ (9a)
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{+}_{n}a\rangle=-i\delta\langle\sigma^{+}_{n}a\rangle-ig_{n}e^{-i\delta t}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+ig_{n}e^{-i\delta t}\langle a^{{\dagger}}a\rangle-F^{*}(t)|\kappa_{n}|^{2}\langle\sigma^{+}_{n}a\rangle,$ (9b)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle=i\delta\langle\sigma^{-}_{n}a^{{\dagger}}\rangle+ig_{n}e^{i\delta t}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-ig_{n}e^{i\delta t}\langle a^{{\dagger}}a\rangle-F(t)|\kappa_{n}|^{2}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle,$ (9c)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle a^{{\dagger}}a\rangle=i\sum_{n}g_{n}\left(e^{i\delta t}\langle\sigma^{+}_{n}a\rangle-e^{-i\delta t}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle\right),$ (9d)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}F(t)=\sum_{n}|\kappa_{n}|^{2}F^{2}(t)-(\gamma+i\Omega-i\omega_{a})F(t)+\frac{\gamma\chi}{2},$ (9e)
@@
where $\chi$ is a constant as in Appendix A. The first two lines of RHS of Eq. (9a) represent that the atomic state excited at the $n$-th energy level can be acquired from a lower energy level by absorbing one photon from the cavity, or from a higher energy level by emitting one photon into the cavity, and the third line of Eq. (9a) represents the non-Markovian interaction between the atom and environment, which can make the atom decay to its ground state. Eq. (9b) and Eq. (9c) represents the interface between the atom and the cavity via the emitting and absorbing processes of a photon. This process is influenced by the detuning between the atom and the cavity (i.e., the first component of the RHS of Eq. (9b) and Eq. (9c)), the coupling strengths $g_{n}$, and the non-Markovian decaying to the environment represented by the last component of the RHS of Eq. (9b) and Eq. (9c), respectively. Eq. (9d) represents how the number of photons in the cavity is influenced by the coupling between the atom and cavity. The nonlinear Eq. (9e) represents the non-Markovian decaying amplitude of the atom to the environment, which is derived in detail in Appendix A.
Additionally, the atom populations are normalized as $\langle\sigma_{00}\rangle+\sum_{n=1}^{N-1}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle=1$, where $\langle\sigma_{00}\rangle$ represents the population that the atom is at the ground state and its dynamics is governed as
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma_{00}\rangle=$ $\left(F(t)+F^{*}(t)\right)\sum_{n=1}^{N-1}|\kappa_{n}|^{2}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$ (10)
$+ig_{1}e^{i\delta t}\langle\sigma^{+}_{1}a\rangle-ig_{1}e^{-i\delta t}\langle\sigma^{-}_{1}a^{{\dagger}}\rangle.$
We define the state vector as
$X(t)=\left[X_{1}(t),X_{2}(t),\cdots,X_{N-1}(t),\langle a^{{\dagger}}a\rangle\right]^{T},$ (11)
@@
where $T$ denotes the matrix transpose and $X_{n}(t)=\left[\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle,\langle\sigma^{+}_{n}a\rangle,\langle\sigma^{-}_{n}a^{{\dagger}}\rangle\right]^{T}$. Then Eq. (9) can be rewritten as
$\frac{\mathrm{d}}{\mathrm{d}t}X$ $=\begin{bmatrix}A_{1}(t)&B_{1}(t)&\cdots&0&D_{1}(t)\\ 0&A_{2}(t)&\cdots&0&D_{2}(t)\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&A_{N-1}(t)&D_{N-1}(t)\\ C_{1}(t)&C_{2}(t)&\cdots&C_{N-1}(t)&0\end{bmatrix}X$ (12)

### v3

@@ I Introduction
Quantum control has garnered much attention due to its potential applications in quantum optics [1, 2], quantum information processing [3], quantum engineering [4] and others [5]. Control methods for quantum systems can be categorized into the open-loop control and closed-loop control, similar to that in classical systems [1]. In open-loop control, designed or iteratively optimized control pulses are utilized to produce required gate operations in quantum computations [6, 7]. When a control field is applied upon an atom, the atom can be excited, leading to the generation of single or multiple photons for quantum networking [8]. On the other hand, closed-loop quantum control can be realized through coherent feedback or measurement feedback methods. For instance, when an atom or cavity quantum electrodynamics (cavity-QED) system is coupled with a waveguide, a coherent feedback channel can be established using photons transmitted in the waveguide [9]. Subsequently, the atomic dynamics and photonic states can be modulated by tuning the parameters of the coherent feedback loop [10, 9, 11]. This form of coherent feedback dynamics can be represented as a linear control system with delays determined by the length of the feedback loop [10, 9]. Then the quantum state dynamics can be interpreted in terms of the stability of a linear control system with delays [12, 11, 13].
@@
In addition to quantum coherent feedback, quantum measurement feedback is another commonly employed method for feedback control in quantum systems [1]. This approach involves designing feedback control based on the measurement outcomes of the quantum system [14, 13]. Within the framework of control theory, quantum measurement feedback control can be represented using stochastic equations due to the noise in measurement and detection apparatuses [15, 16, 17], distinguishing it from feedback control without measurement [18, 15]. By employing these measurement techniques and feedback control, quantum measurement feedback can influence the evolution of quantum states and facilitate the generation of desired quantum states. Consequently, quantum measurement feedback control holds significant promise for various applications in open quantum systems where the quantum states are affected by the environment [1]. Notable applications include the use of measurement-based quantum feedback control in the quantum error correction (QEC) to rectify error bits in quantum computations [19], and preserving the coherence of a quantum state when the quantum system interacts with its environment[20].
@@
In the realm of open quantum system control, it is typically assumed that the environmental evolution timescale is considerably shorter than that of the quantum system of interest, and the decoherence rate of the quantum control system to the environment can be simplified as static [21]. This assumption allows for the modeling of the interaction between the quantum system and the environment using a master equation with a static decaying rate, which is a widely employed Markovian approximation method for studying quantum dynamics in open systems. However, in numerous scenarios, this Markovian approximation proves inadequate for comprehensively analyzing the dynamics of open quantum systems [21, 22]. For instance, in the experimental implementation utilizing nuclear magnetic resonance (NMR) [23], where the NMR qubit interacts with a non-Markovian environment characterized by a randomized configuration of modulated radio-frequency fields. In such setups, the interaction between the NMR and the environment can lead to information backflow from the environment to the qubit, thus the decoherence of the qubit can be nonmonotonic, which is different from the monotonic decoherence of a qubit in the Markovian environment [24]. Moreover, experimental evidence of information backflow induced by non-Markovianity has also been observed in optical systems [25]. Given that traditional quantum control strategies relying on Markovian approximation are not directly applicable in these instances, it is required to explore the open-loop and closed-loop quantum control dynamics for non-Markovian open quantum systems. Additionally, it is worth noting that the traditional Morkovian scenario can be encompassed within the broader framework of non-Markovian settings as a simplified special case.
@@
The interaction between a quantum system and its environment, characterized as non-Markovian, can be effectively represented by the integral stochastic Schrödinger equation [26]. This approach incorporates the influence of the environment on quantum states through a complex-valued stochastic process with an integral kennel relative with historic memory effects in the time domain [27, 26]. Alternatively, after averaging the quantum states over the environmental noise, the non-Markovian dynamics can be described using a master equation featuring time-varying Lindblad components [28, 29], resulting from the memory effect inherent in non-Markovian dynamics [30]. Consequently, various non-Markovian quantum control techniques have been developed including the learning [31, 32] or hybrid control approaches upon quantum systems and the modeled classical non-Markovian environment [33], expanding upon the foundation laid by Markovian quantum control methods [34, 35, 36].
@@
In this paper, we utilize the commonly employed cavity-QED system depicted in Fig. 1 as an example to investigate the dynamics of non-Marovkovian quantum control in a novel manner. We for the first time clarify that, the time-varying parameters indicating a non-Markovian process for open quantum system can be analyzed from the perspective of nonlinear process. Based on this, we illustrate the transition from non-Markovian transient dynamics to steady Markovian dynamics using stability theory in nonlinear control. Subsequently, we describe the evolution of atomic and photonic states in the non-Markovian quantum dynamics using linear time varying (LTV) control equations, where external drives or the measurement feedback can influence the quantum dynamics. Lastly, we extend our analysis to encompass non-Markovian dynamics in systems where atoms in multiple coupled cavities interact with a non-Markovian environment. Then the quantum states can be categorized into the stable and unstable subspaces based on the application of measurement feedback controls.
@@ II Non-Markovian quantum control for open systems
In this section, we introduce the quantum control modeling based on the non-Markovian interactions between a cavity-QED system and its environment. The model is made up of the linear time varying dynamics for the interactions between a multi-level atom and a cavity, and the non-Markovian decaying of the quantum system to the environment governed by nonlinear dynamics.
Figure 1: Quantum control based on non-Markovian interactions between the cavity-QED system and environment.
As illustrated in Fig. 1, a resonant cavity is constructed between two mirrors, one multi-level atom with the energy levels $|0\rangle,|1\rangle,\cdots,|N-1\rangle$ in the format of state vectors is coupled to the cavity. The multi-level atom is simultaneously coupled to the environment modeled as a group of oscillators represented with yellow circles. The cavity can be driven by the field represented with the red arrow, and can be detected or measured via its output channel, which is represented with the blue arrow. Using the measurement information, the feedback operator $G$ can be designed and applied upon the quantum system, thus the closed-loop control with measurement feedback can be realized.
@@
where $\omega_{c}$ is the resonant frequency of the cavity with the annihilation (creation) operator $a$ ($a^{{\dagger}}$), the second component represents a group of oscillators in the environment modeled by the annihilation (creation) operator $b_{\omega}$ ($b_{\omega}^{{\dagger}}$) with different frequencies $\omega$, the third component is the atom Hamiltonian with $\omega_{n}$ represents the energy of the $n$-th level represented with the multiplication of the vector $|n\rangle$ and $\langle n|$, $H_{t}$ is composed with the interaction Hamiltonian between the atom and cavity, as well as the interaction Hamiltonian between the atom and environment.
@@
$\bar{H}_{t}=$ $\sum_{n=1}^{N-1}g_{n}\left(e^{-i\Delta_{n}t}\sigma^{-}_{n}a^{{\dagger}}+e^{i\Delta_{n}t}\sigma^{+}_{n}a\right)$ (2)
$+\sum_{n=1}^{N-1}\sum_{\omega}\chi_{\omega}^{(n)}\left(\sigma^{-}_{n}b_{\omega}^{{\dagger}}+\sigma^{+}_{n}b_{\omega}\right),$
where we denote $\tilde{\omega}_{n}=\omega_{n}-\omega_{n-1}$, $\Delta_{n}=\tilde{\omega}_{n}-\omega_{c}$ is the detuning between the cavity resonant frequency and the transition frequency of two neighborhood atom energy levels, $g_{n}$ represents the coupling strengths between the cavity and different atom energy levels, the lowering operator $\sigma^{-}_{n}=|n-1\rangle\langle n|$, and the rising operator $\sigma^{+}_{n}=|n\rangle\langle n-1|$. The first part on RHS represents the interaction between the atom and cavity, $\sigma^{-}_{n}a^{{\dagger}}$ represents that when the atom decays from the $n$-th energy level to the $(n-1)$-th energy level, it can emit one photon into the cavity. The following Hermitian conjugate $\sigma^{+}_{n}a$ represents that the atom can absorb one photon from the cavity and be excited to a higher energy level. The second part on RHS similarly represents the interaction between the multi-level atom and environment, and the coupling strength between the $n$-th energy level and the environmental oscillator with the frequency $\omega$ is $\chi_{\omega}^{(n)}$. $\sigma^{-}_{n}b_{\omega}^{{\dagger}}$ represents that an emitted filed generated by the atom’s decay from the $n$-th to $(n-1)$-th energy level can be absorbed by the environment, similar for the following Hermitian conjugate. For convenience, we denote $\tilde{g}_{n}=g_{n}e^{i\Delta_{n}t}$ in the following.
The different energy levels of the ladder type atom can be coupled to the environment with the operator [38, 39, 40]
$L=\sum_{n=1}^{N-1}\sqrt{\kappa_{n}}L_{n}=\sum_{n=1}^{N-1}\sqrt{\kappa_{n}}|n-1\rangle\langle n|,$ (3)
where $\kappa_{n}$ represents the decaying rate of the atomic state $|n\rangle$ to the environment and $L_{n}=|n-1\rangle\langle n|$. More details are introduced in Appendix A.
@@
$\frac{\mathrm{d}}{\mathrm{d}t}|\psi(t)\rangle=$ $-i\bar{H}|\psi(t)\rangle$ (4)
$+\left[L|\psi(t)\rangle z_{t}-L^{{\dagger}}\int_{0}^{t}\alpha(t,s)\frac{\delta|\psi(t)\rangle}{\delta z_{s}}\mathrm{d}s\right],$
where $\bar{H}$ represents the interaction component between the atom and cavity in Eq. (2), $z_{t}$ is a complex stochastic process determined by the environment with $E(z_{t}^{*}z_{s})=\alpha(t,s)$, $E(z_{t}z_{s})=0$, and
@@
Generalized from the definition of a quantum Markov process in [43], the quantum dynamics governed by equation (4) is non-Markovian when $\gamma$ is finite, as it includes a memory term wherein the quantum state $|\psi(t)\rangle$ depends on past noise $z_{s}$, weighted by a nonzero kernel $\alpha(t,s)$ for $s<t$ [26].
Alternatively, based on Eq. (4), the quantum dynamics can be equivalently represented by the density matrix $\rho(t)=|\psi(t)\rangle\langle\psi(t)|$, whose dynamics is governed by the following master equation according to the derivation in Appendix B,
$\dot{\rho}=$ $-i[\bar{H},\rho]+\sum_{n=1}^{N-1}\left[F_{n}(t)+F_{n}^{*}(t)\right]L_{n}\rho L_{n}^{{\dagger}}$ (6)
$-\sum_{n=1}^{N-1}\left[F_{n}(t)L_{n}^{{\dagger}}L_{n}\rho+F_{n}^{*}(t)\rho L_{n}^{{\dagger}}L_{n}\right],$
where
$F_{n}(t)=\int_{0}^{t}\alpha(t,s)f_{n}(t,s)\mathrm{d}s,$ (7)
and the integration in Eq. (7) represents the memory effect of the non-Markovian process based on the interaction between the atom and environment.
The reduced quantum Markovian dynamics can be distinguished according to the following remark based on the integral kernel $\alpha(t,s)$.
@@
When $\gamma\rightarrow\infty$, the interaction between the system and the environment is Markovian and $\alpha(t,s)=\delta(t-s)$ [27, 43]. Here,
$F_{n}(t)=\int_{0}^{t}\delta(t-s)f_{n}(t,s)\mathrm{d}s=\frac{1}{2}f_{n}(t,t)\triangleq\frac{\chi_{n}}{2},$ (8)
equals a constant according to Appendix A.
@@ II-A Single multi-level atom coupled with the cavity
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle=-i\tilde{g}_{n}\langle\sigma^{+}_{n}a\rangle+i\tilde{g}_{n}^{*}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle+i\tilde{g}_{n+1}\langle\sigma^{+}_{n+1}a\rangle$
@@
$~{}~{}~{}~{}~{}~{}-i\tilde{g}_{n+1}^{*}\langle\sigma^{-}_{n+1}a^{{\dagger}}\rangle-\left(F_{n}(t)+F_{n}^{*}(t)\right)\kappa_{n}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$
$~{}~{}~{}~{}~{}~{}+\left(F_{n+1}(t)+F_{n+1}^{*}(t)\right)\kappa_{n+1}\langle\sigma^{+}_{n+1}\sigma^{-}_{n+1}\rangle,$ (9a)
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{+}_{n}a\rangle=-i\Delta_{n}\langle\sigma^{+}_{n}a\rangle-i\tilde{g}_{n}^{*}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+i\tilde{g}_{n}^{*}\langle a^{{\dagger}}a\rangle-F_{n}^{*}(t)\kappa_{n}\langle\sigma^{+}_{n}a\rangle,$ (9b)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle=i\Delta_{n}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle+i\tilde{g}_{n}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-i\tilde{g}_{n}\langle a^{{\dagger}}a\rangle-F_{n}(t)\kappa_{n}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle,$ (9c)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle a^{{\dagger}}a\rangle=i\sum_{n}\left(\tilde{g}_{n}\langle\sigma^{+}_{n}a\rangle-\tilde{g}_{n}^{*}\langle\sigma^{-}_{n}a^{{\dagger}}\rangle\right),$ (9d)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}F_{n}(t)=\kappa_{n}F_{n}^{2}(t)-\left(\gamma+i\Omega-i\tilde{\omega}_{n}\right)F_{n}(t)+\frac{\gamma\chi_{n}}{2},$ (9e)
@@
where $\chi_{n}$ is a constant as in Eq. (8) and Appendix A. Eq. (9a) represents that the atomic state excited at the $n$-th energy level can be acquired from a lower energy level by absorbing one photon from the cavity or from a higher energy level by emitting one photon into the cavity, which are represented by the first four components on the RHS, and the following components represent the non-Markovian interaction between the atom and environment, which can make the atom decay to a lower energy level. Eq. (9b) and Eq. (9c) represent the interface between the atom and the cavity via the emitting and absorbing processes of a photon. This process is influenced by the detuning between the atom and the cavity (i.e., the first component of the RHS of Eq. (9b) and Eq. (9c)), the coupling strengths $\tilde{g}_{n}$, and the non-Markovian decaying to the environment represented by the last component of the RHS of Eq. (9b) and Eq. (9c), respectively. Eq. (9d) represents how the number of photons in the cavity is influenced by the coupling between the atom and cavity. The nonlinear Eq. (9e) represents the non-Markovian decaying rate of the atom to the environment, which is derived in detail in Appendix A. Additionally, the atom populations are normalized as $\langle\sigma_{00}\rangle+\sum_{n=1}^{N-1}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle=1$, where $\langle\sigma_{00}\rangle$ represents the population that the atom is at the ground state and its dynamics is governed as
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma_{00}\rangle=$ $\left[F_{1}(t)+F_{1}^{*}(t)\right]\kappa_{1}\langle\sigma^{+}_{1}\sigma^{-}_{1}\rangle$ (10)
$+i\tilde{g}_{1}\langle\sigma^{+}_{1}a\rangle-i\tilde{g}_{1}^{*}\langle\sigma^{-}_{1}a^{{\dagger}}\rangle.$
We define the state vector
$X(t)=\left[X_{1}(t),X_{2}(t),\cdots,X_{N-1}(t),\langle a^{{\dagger}}a\rangle\right]^{\rm T},$ (11)
@@
with $X_{n}(t)=\left[\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle,\langle\sigma^{+}_{n}a\rangle,\langle\sigma^{-}_{n}a^{{\dagger}}\rangle\right]^{\rm T}$ and ${\rm T}$ denoting the matrix transpose. Then Eq. (9) can be rewritten as
$\frac{\mathrm{d}X}{\mathrm{d}t}$ $=\begin{bmatrix}A_{1}(t)&B_{1}(t)&\cdots&0&D_{1}(t)\\ 0&A_{2}(t)&\cdots&0&D_{2}(t)\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&A_{N-1}(t)&D_{N-1}(t)\\ C_{1}(t)&C_{2}(t)&\cdots&C_{N-1}(t)&0\end{bmatrix}X$ (12)